# MARS: Optimizing Dual-System Deep Research via Multi-Agent Reinforcement Learning

### v1

@@ C.3 Reward Design
We introduce a straightforward yet effective reward design for our multi-agent RL training. For each rollout trajectory, we employ LLMs (Qwen2.5-72B-Instruct) as evaluators to assess the correctness of the predicted final answer following HLE’s official prompt [42]. The reward function is defined as:
@@ C.5 Implementation Details
This section provides comprehensive implementation details of our MARS. We initialize our policy model with Qwen2.5-7B-Instruct [56], which serves as the foundation for both System 1 and System 2. Furthermore, we provide external tools such as Google Search, Google Scholar via the SerpAPI [^3] https://serpapi.com/, and Python code interpreter for supporting autonomous tool selection during the reasoning process. We employ the GRPO algorithm [47] with the following hyperparameters: learning rate of 1e-6, batch size of 32, sampled responses per prompt (group size $G$) of 16, temperature of 1.0, top-p of 0.95, and both KL loss coefficient and entropy coefficient set to 0. Additionally, we set different maximum lengths for System 1 and System 2: prompt lengths of 23,552 and 3,072 tokens, and response lengths of 8,192 and 28,672 tokens, respectively. This configuration enables System 1 to better incorporate external knowledge while allowing System 2 to focus on sophisticated multi-step reasoning. For all baselines, we maintain their original settings in their respective papers to ensure optimal performance. Table 5 summarizes the key hyperparameters used during the reinforcement learning phase.
@@
C.6 Baseline Comparisons
For all baseline models, we made every effort to ensure fair comparison by using their original implementations and settings. However, readers may notice that our reported results for WebThinker (QwQ-32B) in Table 1 differ from those in the original paper [29]. This discrepancy stems from two differences:
@@ C.6 Baseline Comparisons
    Evaluation scale: Our evaluation encompasses the complete HLE text-based question set (2,154 questions), providing a more comprehensive assessment than the 500-question sample used in the original WebThinker paper.
@@
    Evaluation criteria: We strictly adhere to HLE’s official evaluation prompt and criteria for all models, including WebThinker. In contrast, the original WebThinker paper employed its own evaluation prompt that resulted in more lenient scoring.
@@ C.7 Dataset Details
  • Single-Hop QA (800 examples, 15.8%): We incorporated 400 randomly sampled examples each from TriviaQA [23] and PopQA [32]. These datasets focus on direct factual knowledge retrieval, with TriviaQA covering a broad range of trivia questions and PopQA specifically targeting popular entities and common knowledge.
  • Multi-Hop QA (1,500 examples, 29.7%): To strengthen the model’s multi-step reasoning capabilities, we included 500 examples each from HotpotQA [58], 2WikiMultihopQA [17], and MuSiQue [51]. These datasets require reasoning across multiple documents or knowledge pieces to arrive at the correct answer.
  • Biology & Medicine (750 examples, 14.9%): To enhance domain-specific knowledge, we sampled 500 examples from PubMedQA [45] and 250 examples from CUPCase [41], covering biomedical research questions and clinical case analysis respectively.
@@ C.8 Experiment Environments
Appendix D Instruction Templates
D.1 Instruction of System 1 & 2 in our MARS
@@
{tool description} When performing a search: 1. **Persistent Actions for Answers**: You can engage in multiple search iterations, delving deeply into the topic to explore all possible aspects until a satisfactory answer is found. 2. **Repeated Verification**: Before presenting a Final Answer, you will **cross-check** and **validate the information** you’ve gathered to confirm its accuracy and reliability. 3. **Attention to Detail**: You will carefully analyze each information source to ensure that all data is current, relevant, and from credible origins. Your reasoning process should be enclosed within <think> </think> tags. If you need external support, make tool calls inside <tool_call> </tool_call> tags. After a tool call, always reassess the result critically and continue your analysis in a new <think> section. Tools are helpful but not always reliable — treat their output with scrutiny. Finally, present your key reasoning and final answer inside <answer> </answer> tags. Do not nest tags. Each tag block must be independent. <User Prompt>
@@
D.2 Instruction for our data curtion pipeline
@@
For the Challenge and Correctness Verification stage (Best-of-16 sampling), we utilized a different approach. Rather than using a standalone evaluation prompt, we leveraged our dual-system framework itself, combining the System 1 and System 2 instructions (detailed in Appendix D) with Google Search integration. This allowed Qwen2.5-72B-Instruct to perform complete reasoning attempts on each question, providing a more authentic assessment of difficulty level and answer verifiability than a static evaluation would allow.
Instruction of Clarity Filtering ### Instruction ### You are a domain expert proficient in various subjects such as Math, Physics, Biology, Humanities, Computer Science, Engineering, and Chemistry. You will be given a question and its corresponding answer. Please identify whether the given QA-pair exhibits "Clarity". Only output 1 for yes and 0 for no. ### Definition ### "Clarity" is defined as the question being clearly and unambiguously stated, and the answer being unique. ### Examples ### {Few-shot Examples}
@@
Instruction of Graduate-level Filtering ### Instruction ### You are a domain expert proficient in various subjects such as Math, Physics, Biology, Humanities, Computer Science, Engineering, and Chemistry. You will be given a question and its corresponding answer. Please identify whether the given QA-pair contains "Expert-level Knowledge." Only output 1 for yes and 0 for no. ### Definition ### "Expert-level Knowledge" is defined as core theories, cutting-edge research, and complex applications studied during advanced undergraduate and graduate levels, typically mastered only by experts or senior researchers in the field. ### Examples ### {Few-shot Examples}
@@
D.3 Instruction for Reward Model
For our reward model, we strictly adhere to the official evaluation criteria from the Humanity’s Last Exam (HLE) benchmark [42].
Instruction of Reward Model (official Instruction from HLE) Judge whether the following [response] to [question] is correct or not based on the precise and unambiguous [correct_answer] below. [question]: {question} [response]: {response} Your judgement must be in the format and criteria specified below: extracted_final_answer: The final exact answer extracted from the [response]. Put the extracted answer as ’None’ if there is no exact, final answer to extract from the response. [correct_answer]: {correct_answer} reasoning: Explain why the extracted_final_answer is correct or incorrect based on [correct_answer], focusing only on if there are meaningful differences between [correct_answer] and the extracted_final_answer. Do not comment on any background to the problem, do not attempt to solve the problem, do not argue for any answer different than [correct_answer], focus only on whether the answers match. correct: Answer ’yes’ if extracted_final_answer matches the [correct_answer] given above, or is within a small margin of error for numerical problems. Answer ’no’ otherwise, i.e. if there if there is any inconsistency, ambiguity, non-equivalency, or if the extracted answer is incorrect. confidence: The extracted confidence score between 0|%| and 100|%| from [response]. Put 100 if there is no confidence score available.

### v2

@@ C.3 Reward Design
We introduce a straightforward yet effective reward design for our multi-agent RL training. For each rollout trajectory, we employ LLMs (Qwen2.5-72B-Instruct) as evaluators to assess the correctness of the predicted final answer following HLE’s official prompt (Phan et al., 2025). The reward function is defined as:
@@ C.5 Implementation Details
This section provides comprehensive implementation details of our MARS. We initialize our policy model with Qwen2.5-7B-Instruct (Yang et al., 2024), which serves as the foundation for both System 1 and System 2. Furthermore, we provide external tools such as Google Search, Google Scholar via the SerpAPI [^3] https://serpapi.com/, and Python code interpreter for supporting autonomous tool selection during the reasoning process. We employ the GRPO algorithm (Shao et al., 2024) with the following hyperparameters: learning rate of 1e-6, batch size of 32, sampled responses per prompt (group size $G$) of 16, temperature of 1.0, top-p of 0.95, and both KL loss coefficient and entropy coefficient set to 0. Additionally, we set different maximum lengths for System 1 and System 2: prompt lengths of 23,552 and 3,072 tokens, and response lengths of 8,192 and 28,672 tokens, respectively. This configuration enables System 1 to better incorporate external knowledge while allowing System 2 to focus on sophisticated multi-step reasoning. For all baselines, we maintain their original settings in their respective papers to ensure optimal performance. Table 7 summarizes the key hyperparameters used during the reinforcement learning phase.
@@
C.6 Baseline Comparisons and Fairness
We provide detailed clarifications on the fairness of our experimental comparisons, addressing potential concerns about model size, training methodology, and tool availability.
Zero RL Setting.
A critical distinction of MARS is that we adopt a “Zero RL” training paradigm: we initialize from the base model (Qwen2.5-7B-Instruct or Qwen3-8B) and train directly with RL without any Supervised Fine-Tuning (SFT). This is a significantly more challenging setting than approaches that use SFT warmup:
  • WebThinker uses SFT on curated trajectories before RL training, providing a strong initialization.
  • C-3PO relies on SFT with a 72B model, leveraging both scale and supervised data.
  • MARS trains from scratch with RL only, yet outperforms these larger, SFT-trained models.
This demonstrates that our multi-agent RL framework, rather than model scale or supervised data quality, is the key driver of performance.
Model Size Comparison.
We use 7B/8B models, significantly smaller than WebThinker (32B) and C-3PO (72B). The fact that MARS outperforms these larger models despite the parameter disadvantage further validates our architectural and algorithmic contributions.
Data Source Parity.
Our training data is curated exclusively from public sources (Appendix B), including MMIQC, WebInstructSub, GeneralThought, and standard QA datasets. We do not use proprietary data or expensive human annotations. This ensures fair comparison with baselines that also use public data sources.
Tool Availability.
MARS uses three tools: Google Search, Google Scholar, and Python Interpreter. Some baselines (e.g., WebThinker) do not use Google Scholar. This difference stems from architectural constraints: baselines without a learned information processor cannot effectively handle the large volume of academic papers that Scholar returns. Our System 1 can efficiently distill this information, enabling effective use of Scholar. Importantly, as shown in our ablation (Section 3.6), the performance gains come from our co-evolution framework, not merely tool availability—the “w/o System 1” setting maintains all tools but shows significant performance degradation.
WebThinker Score Discrepancy.
Readers may notice that our reported WebThinker results differ from the original paper (Li et al., 2025b). This stems from two factors:
@@ C.6 Baseline Comparisons and Fairness
    Evaluation scale: We evaluate on the complete HLE text-based question set (2,154 questions), rather than the 500-question sample used in the original paper.
@@
    Evaluation criteria: We strictly use HLE’s official evaluation prompt for all models. The original WebThinker paper used a custom prompt that resulted in more lenient scoring.
Why Share a Single Checkpoint for System 1 and System 2?
Our design choice to implement both systems within a single LLM (activated through different prompts) is deliberate and motivated by several factors:
  • Alignment with dual-process theory: In cognitive science, System 1 and System 2 are not separate brains but different modes of the same mind. Our single-model design reflects this—one model that can operate in both modes.
  • Synergistic optimization: Sharing parameters enables tighter co-adaptation. System 2 learns to generate purposes that System 1 (sharing the same knowledge) can effectively interpret, and vice versa.
  • Practical efficiency: A two-model design would double memory requirements and inference complexity, limiting practical deployment.
  • Empirical validation: WebThinker’s two-component design (separate reasoning model + fixed extractor) underperforms our unified approach despite using a larger model (32B vs. 8B), suggesting our single-model co-evolution is more effective.
@@ C.7 Dataset Details
  • Single-Hop QA (800 examples, 15.8%): We incorporated 400 randomly sampled examples each from TriviaQA (Joshi et al., 2017) and PopQA (Mallen et al., 2023). These datasets focus on direct factual knowledge retrieval, with TriviaQA covering a broad range of trivia questions and PopQA specifically targeting popular entities and common knowledge.
  • Multi-Hop QA (1,500 examples, 29.7%): To strengthen the model’s multi-step reasoning capabilities, we included 500 examples each from HotpotQA (Yang et al., 2018), 2WikiMultihopQA (Ho et al., 2020), and MuSiQue (Trivedi et al., 2022). These datasets require reasoning across multiple documents or knowledge pieces to arrive at the correct answer.
  • Biology & Medicine (750 examples, 14.9%): To enhance domain-specific knowledge, we sampled 500 examples from PubMedQA (Reese et al., 2024) and 250 examples from CUPCase (Perets et al., 2025), covering biomedical research questions and clinical case analysis respectively.
@@ C.8 Experiment Environments
Appendix D Additional Experimental Results
D.1 Extended RL Training Analysis
(a) Usage Ratio of Python
(b) Usage Ratio of Google Search
(c) Usage Ratio of Google Scholar
(d) Min. Response Len. (Sys. 1)
(e) Mean Response Len.
(f) Max. Response Len. (Sys. 2)
Figure 6: Extended analysis of RL training dynamics. (a-c) Evolution of tool selection preferences. While Google Search emerges as the predominantly chosen tool, the model maintains capability to utilize Python and Google Scholar when appropriate. (d-f) Response length distributions showing that both systems learn to generate more comprehensive outputs as training advances.
Figure 6(a–c) shows how the learned policy allocates tool usage over training. Early in training, the model often relies on a narrow subset of tools (typically Google Search) and under-utilizes specialized tools, reflecting limited exploration and weak credit assignment under trajectory-level rewards. As training progresses, the tool distribution becomes more structured: Google Search remains the default high-recall tool, while Python and Scholar are invoked selectively when the question demands exact computation or academic evidence.
Figure 6(d–f) reports the evolution of response lengths. We observe that System 1’s minimum response length increases over training, suggesting that it learns to consistently produce non-trivial, purpose-conditioned distillations (rather than empty or overly short summaries). Meanwhile, the mean response length increases and System 2’s maximum response length grows, indicating that System 2 learns to sustain longer multi-step reasoning and to integrate evidence across multiple tool calls.
Together, these trends support two conclusions. First, the performance gains are not solely due to calling tools more frequently; instead, training improves the quality of tool use (better timing, better specialization, and better evidence integration). Second, the co-evolution design stabilizes learning under noisy retrieval: System 2 can expand its search when needed, while System 1 limits context dilution by filtering irrelevant content and preserving task-critical entities/constraints.
Appendix E Instruction Templates
E.1 Instruction of System 1 & 2 in our MARS
@@
{tool description} When performing a search: 1. **Persistent Actions for Answers**: You can engage in multiple search iterations, delving deeply into the topic to explore all possible aspects until a satisfactory answer is found. 2. **Repeated Verification**: Before presenting a Final Answer, you will **cross-check** and **validate the information** you’ve gathered to confirm its accuracy and reliability. 3. **Attention to Detail**: You will carefully analyze each information source to ensure that all data is current, relevant, and from credible origins. Your reasoning process should be enclosed within ¡think¿ ¡/think¿ tags. If you need external support, make tool calls inside ¡tool_call¿ ¡/tool_call¿ tags. After a tool call, always reassess the result critically and continue your analysis in a new ¡think¿ section. Tools are helpful but not always reliable — treat their output with scrutiny. Finally, present your key reasoning and final answer inside ¡answer¿ ¡/answer¿ tags. Do not nest tags. Each tag block must be independent. <User Prompt>
@@
E.2 Instruction for our data curtion pipeline
@@
For the Challenge and Correctness Verification stage (Best-of-16 sampling), we utilized a different approach. Rather than using a standalone evaluation prompt, we leveraged our dual-system framework itself, combining the System 1 and System 2 instructions (detailed in Appendix E) with Google Search integration. This allowed Qwen2.5-72B-Instruct to perform complete reasoning attempts on each question, providing a more authentic assessment of difficulty level and answer verifiability than a static evaluation would allow.
Instruction of Clarity Filtering ### Instruction ### You are a domain expert proficient in various subjects such as Math, Physics, Biology, Humanities, Computer Science, Engineering, and Chemistry. You will be given a question and its corresponding answer. Please identify whether the given QA-pair exhibits ”Clarity”. Only output 1 for yes and 0 for no. ### Definition ### ”Clarity” is defined as the question being clearly and unambiguously stated, and the answer being unique. ### Examples ### {Few-shot Examples}
@@
Instruction of Graduate-level Filtering ### Instruction ### You are a domain expert proficient in various subjects such as Math, Physics, Biology, Humanities, Computer Science, Engineering, and Chemistry. You will be given a question and its corresponding answer. Please identify whether the given QA-pair contains ”Expert-level Knowledge.” Only output 1 for yes and 0 for no. ### Definition ### ”Expert-level Knowledge” is defined as core theories, cutting-edge research, and complex applications studied during advanced undergraduate and graduate levels, typically mastered only by experts or senior researchers in the field. ### Examples ### {Few-shot Examples}
@@
E.3 Instruction for Reward Model
For our reward model, we strictly adhere to the official evaluation criteria from the Humanity’s Last Exam (HLE) benchmark (Phan et al., 2025).
Instruction of Reward Model (official Instruction from HLE) Judge whether the following [response] to [question] is correct or not based on the precise and unambiguous [correct_answer] below. [question]: {question} [response]: {response} Your judgement must be in the format and criteria specified below: extracted_final_answer: The final exact answer extracted from the [response]. Put the extracted answer as ’None’ if there is no exact, final answer to extract from the response. [correct_answer]: {correct_answer} reasoning: Explain why the extracted_final_answer is correct or incorrect based on [correct_answer], focusing only on if there are meaningful differences between [correct_answer] and the extracted_final_answer. Do not comment on any background to the problem, do not attempt to solve the problem, do not argue for any answer different than [correct_answer], focus only on whether the answers match. correct: Answer ’yes’ if extracted_final_answer matches the [correct_answer] given above, or is within a small margin of error for numerical problems. Answer ’no’ otherwise, i.e. if there if there is any inconsistency, ambiguity, non-equivalency, or if the extracted answer is incorrect. confidence: The extracted confidence score between 0—%— and 100—%— from [response]. Put 100 if there is no confidence score available.